# Causal-Informed Contrastive Learning: Towards Bias-Resilient Pre-training under Concept Drift

### v1

@@ B.1 Datasets
ImageNet-LT [10] is a long-tailed subset derived from the ImageNet [9] dataset, designed to benchmark machine learning models under realistic class imbalance. It contains 115k training images across 1,000 categories, with class frequencies ranging from 1,280 (head classes) to merely 5 (tailed classes), simulating real-world data skewness. The validation set is balanced (20 images per class), while the test set aligns with the standard ImageNet validation data (50,000 images). It serves as a pivotal resource for studying long-tail recognition challenges, including few-shot learning, open-set generalization, and algorithmic fairness, while retaining compatibility with the original ImageNet hierarchy and evaluation protocols.
@@
iNaturalist2018 [29] is a large-scale dataset designed to study fine-grained recognition under naturally occurring long-tailed class distributions. It comprises 675k training and validation images spanning 5,089 species hierarchically organized into 13 super-categories, with severe imbalance reflecting ecological rarity. The dataset integrates taxonomic metadata, geospatial coordinates, and temporal annotations, supporting research in hierarchical learning, transfer learning, and self-supervised pretraining.
@@
In terms of OOD detection, the Texture (DTD) [43] is a comprehensive collection of 5,640 texture images, meticulously organized into 47 categories based on human-centric perceptual attributes. Designed to support texture analysis and recognition tasks, DTD is divided into training, validation, and test sets, each containing 40 images per category. Besides, the iNat-OOD dataset [45] is a specialized subset of the iNaturalist [29] dataset, designed for evaluating out-of-distribution (OOD) detection methods in large-scale image classification tasks. The dataset is particularly valuable for testing the robustness of models in identifying novel or unseen categories, especially in ecological and biodiversity contexts. Meanwhile, ImageNet-O [46] is a specialized dataset designed to evaluate the robustness of visual models in detecting out-of-distribution (OOD) samples. It consists of 2,000 images from classes not included in the standard ImageNet-1k [9] dataset, making it a valuable benchmark for testing OOD detection methods. Additionally, OpenImage-O [47]is a large-scale, manually annotated dataset designed for out-of-distribution (OOD) detection tasks. It is derived from the OpenImage-V3 test set, which contains a diverse and natural distribution of images collected from Flickr without predefined class names or tags. This dataset aims to overcome the limitations of existing OOD benchmarks by providing a more realistic and challenging testbed for evaluating the robustness of computer vision models
@@
Regarding domain shift, ImageNet-V2 [40] is a comprehensive test set designed to evaluate the robustness and generalization of image classification models. It comprises 10,000 images, with 10 images per class, closely following the original labeling protocol of ImageNet. This dataset is instrumental in assessing how well models trained on the original ImageNet dataset can generalize to new, unseen data, making it a valuable resource for advancing computer vision research. The ImageNet-Sketch [41] is a unique collection of 50,000 hand-drawn sketch images, with 50 images for each of the 1,000 ImageNet classes. Constructed using Google Image queries, this dataset is designed to evaluate models’ ability to learn out-of-domain semantics at the ImageNet scale. ImageNet-R [42] is a collection of 30k images representing 200 ImageNet [9] classes. These images are artistic renditions, including art, cartoons, graffiti, embroidery, origami, paintings, and other forms of creative expressions. The dataset is designed to test the robustness and generalization capabilities of image classification models when faced with non-standard inputs .

### v2

@@ B.1 Datasets
ImageNet-LT (liuLargeScaleLongTailedRecognition2019) is a long-tailed subset derived from the ImageNet (russakovskyImagenetLargeScale2015) dataset, designed to benchmark machine learning models under realistic class imbalance. It contains 115k training images across 1,000 categories, with class frequencies ranging from 1,280 (head classes) to merely 5 (tailed classes), simulating real-world data skewness. The validation set is balanced (20 images per class), while the test set aligns with the standard ImageNet validation data (50,000 images). It serves as a pivotal resource for studying long-tail recognition challenges, including few-shot learning, open-set generalization, and algorithmic fairness, while retaining compatibility with the original ImageNet hierarchy and evaluation protocols.
@@
iNaturalist2018 (vanhornINaturalistSpeciesClassification2018a) is a large-scale dataset designed to study fine-grained recognition under naturally occurring long-tailed class distributions. It comprises 675k training and validation images spanning 5,089 species hierarchically organized into 13 super-categories, with severe imbalance reflecting ecological rarity. The dataset integrates taxonomic metadata, geospatial coordinates, and temporal annotations, supporting research in hierarchical learning, transfer learning, and self-supervised pretraining.
@@
In terms of OOD detection, the Texture (DTD) (cimpoi2014describing) is a comprehensive collection of 5,640 texture images, meticulously organized into 47 categories based on human-centric perceptual attributes. Designed to support texture analysis and recognition tasks, DTD is divided into training, validation, and test sets, each containing 40 images per category. Besides, the iNat-OOD dataset (huangMOSScalingOutofDistribution2021) is a specialized subset of the iNaturalist (vanhornINaturalistSpeciesClassification2018a) dataset, designed for evaluating out-of-distribution (OOD) detection methods in large-scale image classification tasks. The dataset is particularly valuable for testing the robustness of models in identifying novel or unseen categories, especially in ecological and biodiversity contexts. Meanwhile, ImageNet-O (hendrycksNaturalAdversarialExamples2021) is a specialized dataset designed to evaluate the robustness of visual models in detecting out-of-distribution (OOD) samples. It consists of 2,000 images from classes not included in the standard ImageNet-1k (russakovskyImagenetLargeScale2015) dataset, making it a valuable benchmark for testing OOD detection methods. Additionally, OpenImage-O (wangViMOutofDistributionVirtualLogit2022)is a large-scale, manually annotated dataset designed for out-of-distribution (OOD) detection tasks. It is derived from the OpenImage-V3 test set, which contains a diverse and natural distribution of images collected from Flickr without predefined class names or tags. This dataset aims to overcome the limitations of existing OOD benchmarks by providing a more realistic and challenging testbed for evaluating the robustness of computer vision models
@@
Regarding domain shift, ImageNet-V2 (rechtImageNetClassifiersGeneralize2019) is a comprehensive test set designed to evaluate the robustness and generalization of image classification models. It comprises 10,000 images, with 10 images per class, closely following the original labeling protocol of ImageNet. This dataset is instrumental in assessing how well models trained on the original ImageNet dataset can generalize to new, unseen data, making it a valuable resource for advancing computer vision research. The ImageNet-Sketch (wangLearningRobustGlobal2019) is a unique collection of 50,000 hand-drawn sketch images, with 50 images for each of the 1,000 ImageNet classes. Constructed using Google Image queries, this dataset is designed to evaluate models’ ability to learn out-of-domain semantics at the ImageNet scale. ImageNet-R (hendrycksManyFacesRobustness2021) is a collection of 30k images representing 200 ImageNet (russakovskyImagenetLargeScale2015) classes. These images are artistic renditions, including art, cartoons, graffiti, embroidery, origami, paintings, and other forms of creative expressions. The dataset is designed to test the robustness and generalization capabilities of image classification models when faced with non-standard inputs .